# Data-Driven Chance-Constrained Multiple-Choice Knapsack Problem: Model, Algorithms, and Applications

### v1

@@ Abstract
The multiple-choice knapsack problem (MCKP) is a classic NP-hard combinatorial optimization problem. Motivated by several significant practical applications, this work investigates a novel variant of MCKP called data-driven chance-constrained multiple-choice knapsack problem (DDCCMCKP), where the item weight is a random variable with unknown probability distribution. We first present the problem formulation of DDCCMCKP, and then establish two benchmark sets. The first set contains synthetic instances, and the second set is devised to simulate a real-world application scenario of a certain telecommunication company. To solve DDCCMCKP, we propose a data-driven adaptive local search (DDALS) algorithm. The main merit of DDALS lies in evaluating solutions with chance constraints by data-driven methods, under the condition of unknown distributions and only historical sample data being available. The experimental results demonstrate the effectiveness of the proposed algorithm and show that it is superior to other baselines. Additionally, ablation experiments confirm the necessity of each component in the algorithm. Our proposed algorithm can serve as the baseline for future research, and the code and benchmark sets will be open-sourced to further promote research on this challenging problem.
@@
Combinatorial optimization, data-driven approach, multiple-choice knapsack problem, evolutionary algorithm.
@@ I Introduction
The multiple-choice knapsack problem (MCKP) is a variant of the classic knapsack problem [1]. MCKP finds practical applications in various areas, such as supply chain optimization [2], target market selection [3], investments portfolio selection [4], transportation programming [5], and component selection in the information technology systems [6]. Specifically, MCKP involves a set of items, each possessing a weight and a value. The items are divided into several disjoint classes and only one item can be selected in each class [1, 7]. The objective of MCKP is to select exactly one item from each class such as to maximize the total value while adhering to a predefined weight limit (i.e., knapsack constraint). Over the past several decades, plenty of methods for solving MCKP have been proposed, including greedy methods [8], branch-and-bound [9], dynamic programming [10], approximation algorithms [11], and heuristic search [12].
Despite of significant research success in addressing the MCKP, those advanced methods are incapable of handling problems in real-world scenarios where item weights are randomly distributed instead of fixed. To handle random variables, a common approach is to model deterministic constraints as chance constraints, i.e., the probability (confidence level) of total weight (random variable) of the selected items not exceeding a given deterministic value is no less than a certain threshold [13, 14]. Furthermore, the specific probability distributions of the random variables could be unknown [15, 16]. Instead, one only has historical sample data to characterize the random variables. The collection of sample data for each item is assumed to be independent and uncorrelated. This variant of MCKP is referred to as data-driven chance-constrained multiple-choice knapsack problem (DDCCMCKP).
@@
A typical application of DDCCMCKP is the configuration of fifth-generation (5G) end-to-end (E2E) network. In the area of 5G for business [17], user plane time delay is one of the key performance indicators for low latency communication [18, 19]. As illustrated in Figure 1, the complete packet transmission path in a 5G cellular network consists of several segments, i.e., eNB, backhaul, core network, Internet/cloud, and the transport between core network and Internet/cloud. More specifically, these segments correspond to classes, where for each class there exist different options (items) and only one option can be selected to implement the class. The total time delay of a packet transmission can be calculated as the sum of the delays induced by all the selected options, i.e., $(T_{eNB}+T_{Backhaul}+T_{Core}+T_{Transport}+T_{Internet})*2$. Moreover, the packet transmission delays of the implementation options are random variables with unknown distributions as they are affected by various factors, including workloads, network configurations, broadcast environment, and random failures of software and hardware. The configuration of 5G E2E network is to select an implementation option for each segment. Different configurations have different total costs and time delays, where the costs are fixed values, but time delays are unknown random variables. The objective is to minimize the total cost of network configuration while ensuring that the probability of meeting the delay is less than a given threshold under a certain value.
@@
Another application of DDCCMCKP is supply chain planning and optimization in modern industries [2, 3]. The supply-chain operations reference (SCOR) model, published by the Supply-Chain Council [20], is a useful tool for the upper management of a firm to design and reconfigure the supply chain to achieve better performance. In SCOR, the E2E supply chain consists of four distinct processes: source, make, deliver, and plan (see Figure 2). Specifically, these processes can be further subdivided into supplier selection, R&D testing, production, transportation, and recycling handling, which correspond to different classes along the supply chain. Therefore, each class has various configuration options that can be influenced by random factors, such as market demand, cost fluctuations, and technology configurations.
@@
Since DDCCMCKP has not been paid research attention, it merits thorough investigation due to its extensive and noteworthy applications. Compared to MCKP, DDCCMCKP faces greater challenges due to the complexity of its constraints. There are two main difficulties: it is a chance-constrained optimization problem, and the distributions of the random variables are unknown. Therefore, a data-driven approach that does not rely on specific distribution information but rather available sample data to evaluate solutions must be proposed.
The main contributions are summarized as follows:
@@
    From the problem perspective, we define DDCCMCKP from several real-world scenarios and formulate the model. To the best of our knowledge, this is the first endeavor to address this specific problem. To further promote research on this problem, we have designed two sets of benchmarks. The first set is a simulated benchmark with a simple mechanism, while the other set is to simulate a real-world business requirement of a telecommunication company. The generated benchmarks will be made publicly available to promote further research.
@@
    From the algorithmic perspective, we propose a novel algorithm named data-driven adaptive local Search (DDALS). The first novelty of DDALS is that we evaluate solutions with chance constraints by a data-driven method. The second is that we design surrogate weights using statistics of sample data to ensure the performance of the search algorithm. Furthermore, our algorithm can adaptively jump between small and larger search regions for efficient searching of high-quality solutions. Lastly, we improve the proportion of feasible solutions through carefully designed filtering mechanisms.
@@
    From the computational perspective, we experimentally confirm the effectiveness of DDALS. We adapt several commonly-used approaches as the baselines, including greedy search (Greedy), genetic algorithm (GA), and estimation of distribution algorithm (EDA). Experimental results show that DDALS outperforms other baseline methods. The proposed algorithm can be served as a baseline for comparison experiments on this challenging problem.
The rest of the paper is organized as follows. In Section II, we present a comprehensive review of related works. Section III provides a formal definition of DDCCMCKP. In Section IV, DDALS is introduced and a detailed explanation of its implementation is provided. In Section V, the benchmarks we designed are described in details. The experimental results and analyses are given in Section VI. Finally, we summarize our findings and provide suggestions for future research in Section VII.
@@ II Related works
In this section, we will review the relevant literature. We first review existing research on the traditional MCKP. The second category of research focuses on chance-constrained optimization. Since the methods developed in robust optimization can solve chance-constrained optimization by constructing proper uncertainty sets [21, 22], we further review robust optimization and distributionally robust optimization. While robust optimization approach minimizes the worst-case scenario under the uncertainty of random variables, the goal of distributionally robust optimization is to ensure the robustness of the optimal solution in the distribution-unknown situations.
@@ II-A Algorithms for MCKP
There have been various stochastic variations of the classical 0-1 MCKP that have been extensively studied in the literature. For instance, Dyer et al.[9] proposed a branch and bound method to solve the MCKP. Pisinger [23] presented a minimal algorithm including linear relaxation, greedy algorithm and dynamic programming. Lamanna et al. [24] introduced a new algorithm for solving the multidimensional MCKP based on Kernel search, which divided the solution process of the algorithm into two phases with complementary objectives. Gens and Levner [11] proposed a fast approximation algorithm for the MCKP with a 4/5 approximation ratio of the optimal solution using binary search. He et al. [25] developed an improved approximate binary search algorithm with an approximation ratio of $3+(1/2)^{t}$. Hifi et al. [12] proposed a reactive local search algorithm for the multidimensional MCKP. Mkaouar et al. [26] introduced a method based on the artificial bee colony algorithm to solve the multidimensional MCKP. In general, existing researches on MCKP cannot be applied, since the solutions in DDCCMCKP can not be evaluated by deterministic approaches.
@@ II-B Chance-Constrained Optimization
Chance-constrained optimization plays a critical role in stochastic optimization [27, 13, 14]. Kleinberg et al. [28] focused on packing multiple bursty connections with varying transmission rates on a link and considered stochastic knapsack problems with known Bernoulli distribution. Goel and Indyk [29] studied the knapsack problem when the jobs had stochastic processing requirements or sizes. They proposed polynomial time approximation scheme (PTAS) when the jobs were all known to follow either Poisson or exponential distributions. Goyal and Ravi[30] proposed a PTAS that used a parametric linear programming reformulation and yielded near-optimal profit efficiently. The random variables were assumed to follow a known Gaussian distribution. Shabtai et al. [31] concentrated on the chance-constrained knapsack problem with known normal distributions. They presented a fast PTAS for the problem that run in poly($n/\epsilon$) time. Zhang et al. [32] studied trajectory prediction with chance constraint, assuming that future position of the moving obstacle could be represented as Gaussian distributions based on the predicted statistics. Chai et al. [33] studied trajectory optimization with chance constraint with known probability density function of random variables defined on the measurable set. Yang and Chakraborty [34] exploited the geometry of the non-convex feasible region of the chance-constrained knapsack problem in the variance-mean plane and transformed the problem into a sequence of deterministic knapsack problems to find an optimal solution. All these works mainly focused on theoretical analysis of situations that the distribution of random variables is known, which are not applicable to our problem as well.
II-C Robust Optimization
A robust optimization approach aims to construct an uncertainty set $\mathcal{U}_{\epsilon}\subseteq\mathcal{U}$ such that a feasible solution of the form $g(x,\xi)\leq 0,\forall\xi\in\mathcal{U}_{\epsilon}$ will also be feasible with probability at least $1-\epsilon$ with respect to $P$, i.e., $\mathcal{P}(g(x,\xi)\leq 0)\geq 1-\epsilon,\forall\xi\in\mathcal{U}$, which has the same form with chance constraint. Han [21] considered a chance-constrained knapsack problem where each item had an independent normally distributed random weight with a known mean and variance, and proposed an efficient pseudo-polynomial time algorithm based on the robust optimization approach. Joung and Lee[22] proposed a robust optimization-based heuristic algorithm where the optimal solution could be obtained by solving the ordinary binary knapsack problem iteratively.
Distributionally robust optimization was proposed as an improvement of robust optimization when the distribution of the random variable was unknown and only observable through a finite training dataset. Mohajerin Esfahani and Kuhn [35] showed that distributionally robust optimization problems over Wasserstein balls could be reformulated as finite convex programs. Ji and Lejeune [15] studied distributionally robust chance-constrained programming optimization problems with data-driven Wasserstein ambiguity sets. The proposed algorithmic and reformulation framework were applied to all types of distributionally robust chance-constrained optimization problems, including the multidimensional knapsack problem. Additionally, Xie [16] proposed that the distributionally robust chance-constrained problem could be reformulated as a conditional value-at-risk constrained optimization problem. The problem with binary decision variables admitted a big-M free formulation and could be solved by a branch and cut algorithm. Gao and Kelywegt [36] derived a strong duality reformulation of the corresponding distributionally robust stochastic optimization problem and construct approximate worst-case distributions explicitly. Note that most of these work can not be applied. We have tried to make modifications to Ji’s [15] and Yang’s [34] algorithm to make it applicable to our problem, but the results show that both of them exhibited unsatisfactory performance. (see Section VI-F for details).
@@ III Notations and problem definition
We give the formulation of DDCCMCKP in Eq. III. Consider $m$ mutually disjoint classes $N_{1},\ldots,N_{m}$ of items. Each item $j\in N_{i}$ has an associated cost $c_{ij}$ and weight $w_{ij}$, where $c_{ij},j\in N_{i},\forall i\in\mathcal{M}=\{1,\ldots,m\}$ are fixed values while $w_{ij},j\in N_{i},\forall i\in\mathcal{M}$ are unknown random variables. The weight of item $j\in N_{i}$ is characterized by $L$ sample data, denoted as $d_{ijl},i\in\{1,\ldots,m\},j\in N_{i},l\in\{1,\ldots,L\}$. Note that the sample data of each item is independent and identically distributed (i.i.d.). The goal is to select one item from each class to minimize the total cost, subject to keeping the total weight below the knapsack’s capacity $W$ with probability higher than the given confidence level $P_{0}$. By introducing the binary variables $x_{ij}$ which take on value 1 if the item $j$ is chosen in class $N_{i}$, the problem can be formulated as:
@@
To avoid trivial solutions, we assume that the following inequality holds:
@@
Note that MCKP is NP-hard [1]. Supposing that all samples of each item’s weight $w_{ij}$ are the same, then DDCCMCKP reduces to traditional MCKP. Therefore, MCKP can be seen as a special case of DDCCMCKP, which indicates that DDCCMCKP is also NP-hard. It motivates us to adopt an algorithm based on a heuristic search framework to solve DDCCMCKP.
@@ IV Data-driven adaptive local search
As mentioned in Section II, although there have been many works in the fields of MCKP, chance-constrained optimization and robust optimization, none of the methods could be directly applied to our problem. In order to tackle the DDCCMCKP, we present a novel approach called data-driven adaptive local search (DDALS) in this work. Specifically, DDALS consists of three main parts, confidence level evaluation, surrogate weight design and local search framework. Finally, we will analyze the complexity of the algorithm.
@@ IV-A Confidence Level Evaluation
Since the sample data is assumed to be i.i.d., it is necessary to make full use of the information contained in the sample data, since we have no information about the true distribution. Given a solution $S$, a natural idea is to calculate all possible combinations of the data and count the proportion of these combinations that meet the requirements. This proportion is an estimate of the real confidence level $P$ of the solution. However, obviously, a brute-force method is not applicable since it requires a calculation complexity of $L^{m}$, which is related to the problem size $m$ and sample size $L$. It may be feasible for small-scale problems, but it becomes impractical for large-scale problems. Therefore, based on the above considerations, we propose an exact evaluation method for small-scale problems that is much faster than brute-force enumeration. For large-scale problems, we propose an accelerated Monte Carlo evaluation method.

### v2

# Chance-Constrained Multiple-Choice Knapsack Problem: Model, Algorithms, and Applications

@@ Abstract
The multiple-choice knapsack problem (MCKP) is a classic NP-hard combinatorial optimization problem. Motivated by several significant real-world applications, this work investigates a novel variant of MCKP called chance-constrained multiple-choice knapsack problem (CCMCKP), where the item weights are random variables. In particular, we focus on the practical scenario of CCMCKP, where the probability distributions of random weights are unknown but only sample data is available. We first present the problem formulation of CCMCKP and then establish two benchmark sets. The first set contains synthetic instances and the second set is devised to simulate a real-world application scenario of a certain telecommunication company. To solve CCMCKP, we propose a data-driven adaptive local search (DDALS) algorithm. The main novelty of DDALS lies in its data-driven solution evaluation approach that can effectively handle unknown probability distributions of item weights. Moreover, in cases with unknown distributions, high intensity of chance constraints, limited amount of sample data and large-scale problems, it still exhibits good performance. Experimental results demonstrate the superiority of DDALS over other baselines on the two benchmarks. Additionally, ablation studies confirm the effectiveness of each component of the algorithm. Finally, DDALS can serve as the baseline for future research, and the benchmark sets are open-sourced to further promote research on this challenging problem.
@@
Combinatorial optimization, data-driven optimization, chance-constrained multiple-choice knapsack problem, evolutionary algorithm.
@@ I Introduction
The multiple-choice knapsack problem (MCKP) is a variant of the classic knapsack problem [1]. MCKP finds applications in various areas, such as supply chain optimization [2], target market selection [3], investments portfolio selection [4], transportation programming [5], and component selection in the information technology systems [6]. Specifically, MCKP involves a set of items, each possessing a weight and a value. The items are divided into several disjoint classes. The objective of MCKP is to select exactly one item from each class [1, 7] such as to maximize the total value while adhering to a predefined weight limit (i.e., the knapsack constraint). Over the past several decades, plenty of methods for solving MCKP have been proposed, including greedy methods [8], branch-and-bound [9], dynamic programming [10], approximation algorithms [11], and heuristic search [12].
Despite the research progress made thus far, advanced methods for solving MCKP still fall short when confronted with scenarios where item weights are not fixed but are random variables. In such cases, the constraint is no longer the total weight not exceeding a fixed limit, but rather the probability (confidence level) of the total weight not exceeding the limit is no less than a certain threshold. In this work, we refer to this particular variant of MCKP as the chance-constrained multiple-choice knapsack problem (CCMCKP). Furthermore, in real-world applications of CCMCKP, the probability distributions of item weights are often unknown [13, 14], and sample data is the only available information for characterizing these random variables.
@@
The above-described CCMCKP finds applications in various domains. A typical example is the configuration of fifth-generation (5G) end-to-end (E2E) network. In the area of 5G for business [15], user plane time delay is one of the key performance indicators for low latency communication [16, 17]. As illustrated in Figure 1, the complete packet transmission path in a 5G cellular network consists of several segments, such as eNB, backhaul, and core network. Specifically, these segments correspond to the classes in CCMCKP, where for each class there exist different implementation options (items) and only one option can be selected. The total time delay of a packet transmission is calculated as the sum of the delays induced by all the selected options. Here, packet transmission delays (item weights) are random variables with unknown probability distributions because they are generally affected by various unpredictable factors such as workloads, broadcast environment, as well as random failures of software and hardware. The configuration of 5G E2E network is to select an implementation option for each segment, such that the total cost of the network configuration is minimized while ensuring the probability of packet transmission delay not exceeding the limit is no less than a certain threshold.
@@
Another application of CCMCKP is supply chain planning and optimization in modern industries [2, 3]. The supply-chain operations reference (SCOR) model, published by the Supply-Chain Council [18], is a useful tool to design and reconfigure the supply chain. As illustrated in Figure 2, in SCOR, the E2E supply chain consists of four distinct processes: source, make, deliver, and plan. Specifically, these processes can be further subdivided into supplier selection, R&D testing, production, transportation, and recycling handling, corresponding to the classes in CCMCKP. Each class has different configuration options and only one option can be selected. The processing delays of these options are random variables with unknown probability distributions because they are influenced by various random factors, such as market demand, cost fluctuations, and technology configurations. Finally, the goal is to select an option for each class to minimize the total cost, while ensuring the probability of total processing delay not exceeding the limit is no less than a certain threshold.
Despite the practical significance of CCMCKP, currently there is no direct research on solving this challenging problem. On the other hand, there are alternative methods that can be adapted to solve CCMCKP, such as stochastic optimization (SO) [19] that assumes known probability distributions of item weights, and distributionally robust optimization (DRO) [13] that employs the uncertainty set technique. Nonetheless, these methods exhibit unsatisfactory performance when the assumed probability distribution does not hold, intensity of the chance constraint is high, or the amount of sample data is not sufficiently large (see Section VI-B). Motivated by these observations, we propose a novel data-driven adaptive local search (DDALS) algorithm to effectively address CCMCKP. Through rigorous experiments, we demonstrate the superior performance of DDALS in comparison to existing methods.
@@
The main contributions of this work are summarized below.
@@
    From the optimization problem perspective, we formulate the model of CCMCKP. To the best of our knowledge, this is the first endeavor to address this specific problem. Then, we establish two benchmark sets of CCMCKP. The first set contains synthetic instances while the second set is devised to simulate a real-world business requirement of a certain telecommunication company. The generated benchmarks are available online to promote further research.
@@
    From the algorithmic perspective, we propose DDALS to solve CCMCKP. The main novelty of DDALS lies in its data-driven solution evaluation approach that can effectively handle situations where the distributions of item weights are unknown. Besides, DDALS incorporate surrogate weights using statistics of sample data and can adaptively jump between small and large search regions to find high-quality solutions. Lastly, DDALS employs a carefully designed mechanism to filter out infeasible solutions, further enhancing its efficiency.
@@
    From the computational perspective, we experimentally confirm the effectiveness of DDALS. First, experimental results show the superiority of DDALS over SO and DRO methods in cases with unknown distributions, high intensity of chance constraints, limited amount of sample data and large-scale problems. Then, further experiments show that DDALS outperforms alternative commonly used optimization methods including greedy search (Greedy), genetic algorithm (GA), and estimation of distribution algorithm (EDA). Lastly, ablation studies confirm the effectiveness of each component of DDALS. For future research, DDALS can serve as a baseline method for comparison experiments on solving CCMCKP.
The rest of the paper is organized as follows. Section II presents the review of related works. Section III provides a formal definition of CCMCKP. In Section IV, DDALS is introduced and a detailed explanation is provided. In Section V, the benchmarks are described in details. Experimental results and analyses are given in Section VI. Finally, we summarize our findings and provide suggestions for future research in Section VII.
@@ II Related works
In this section, we first review existing research on the classic MCKP and then introduce the research on chance-constrained optimization. Finally, we review DRO that solves chance-constrained optimization problems by constructing uncertainty sets.
@@ II-A Algorithms for MCKP
Classic MCKP has been extensively studied in the literature. For instance, Dyer et al. [9] proposed a branch and bound method to solve MCKP. Pisinger [20] presented a minimal algorithm including linear relaxation, greedy algorithm and dynamic programming. Gens and Levner [11] and He et al. [21] both proposed approximation algorithms for MCKP. Hifi et al. [12] presented a reactive local search algorithm, Mkaouar et al. [22] employed the artificial bee colony algorithm, and Lamanna et al. [23] introduced a Kernel Search-based approach for solving multidimensional MCKP. However, these deterministic methods cannot be applied to CCMCKP as they cannot handle stochastic item weights.
@@ II-B Chance-Constrained Optimization
Chance-constrained optimization plays a critical role in SO [24, 25, 26]. Prior works have investigated chance-constrained optimization problems with known probability distributions, including Bernoulli, Poisson, exponential, and Gaussian [27, 28, 29, 30, 31]. Yang and Chakraborty [19] solved the chance-constrained knapsack problem where item weights follow Gaussian distributions. We sought to adapt their method, referred to as the Gaussian-distribution stochastic optimization (GDSO), to address CCMCKP. Specifically, a statistical fitting approach is applied to sample data to obtain the mean and variance of the Gaussian distribution, based on which GDSO is run. However, its performance on solving CCMCKP is not satisfactory due to deviations between the assumed Gaussian distributions and the unknown underlying distributions. Besides, as the problem scale increases, GDSO’s runtime becomes prohibitively long, as the algorithm utilizes dynamic programming. The detailed results are presented in Section VI-B.
II-C Distributionally Robust Optimization (DRO)
Robust optimization aims to construct an uncertainty set $\mathcal{U}_{\epsilon}\subseteq\mathcal{U}$ such that a feasible solution of the form $g(x,\xi)\leq 0,\forall\xi\in\mathcal{U}_{\epsilon}$ is also feasible with probability at least $1-\epsilon$ with respect to $P$, i.e., $\mathcal{P}(g(x,\xi)\leq 0)\geq 1-\epsilon,\forall\xi\in\mathcal{U}$, which has the same form as the chance constraint. DRO is proposed as an improvement of robust optimization when the distributions of the random variables are unknown and only observable through sample data. Researchers have explored various reformulations and approaches for DRO, utilizing techniques such as Wasserstein balls, conditional value-at-risk constraints, and strong duality formulations [32, 14, 13, 33]. Ji and Lejeune [13] studied distributionally robust multidimensional knapsack problem with Wasserstein ambiguity sets. We adapted their method, denoted as DRO, to address CCMCKP. However, experimental results show that DRO exhibits poor performance, when the sample size is not sufficiently large, or when intensity of chance constraints is high. Similar to GDSO, DRO also experiences a significant increase in runtime as the problem scale grows, eventually becoming impractical. The detailed results are presented in Section VI-B.
@@ III Notations and problem definition
We give the formulation of CCMCKP in Eq. III. Consider $m$ mutually disjoint classes $N_{1},\ldots,N_{m}$ of items. Each item $j\in N_{i}$ has an associated cost $c_{ij}$ and weight $w_{ij}$, where $c_{ij},j\in N_{i},\forall i\in\mathcal{M}=\{1,\ldots,m\}$ are fixed values while $w_{ij}$ are random variables. The goal is to select one item from each class to minimize the total cost, subject to keeping the total weight below the knapsack’s capacity $W$ with probability higher than the given confidence level $P_{0}$. By introducing the binary variables $x_{ij}$ which take on value 1 if item $j$ is chosen in class $N_{i}$, the problem is formulated as:
@@
In this work, we focus on the practical scenario where the probability distributions of random variables are unknown. For each item, $L$ sample data is collected, denoted as $d_{ijl},i\in{1,\ldots,m},j\in N_{i},l\in{1,\ldots,L}$, which is assumed to be independent and identically distributed (i.i.d.). To avoid trivial solutions, we also assume that the following inequality holds:
@@
Note that MCKP is proven to be NP-hard [1]. Consider the scenario where item’s weights $w_{ij}$ follow a particular distribution which takes probability 1 at one value, then CCMCKP simplifies to classic MCKP. Consequently, MCKP can be regarded as a special case of CCMCKP, implying that CCMCKP is also NP-hard.
@@ IV Data-driven adaptive local search
As outlined in Section II, currently there is no direct research on solving CCMCKP. To address this problem, a novel approach termed DDALS is introduced. We elaborate three core components of DDALS sequentially: confidence level evaluation, surrogate weight design, and local search framework. Finally, a detailed analysis of the computational complexity is conducted.
@@ IV-A Confidence Level Evaluation
As the sample data is assumed to be i.i.d., it is imperative to extract the full informational value from the sample data. Given a solution $S$, one approach is to calculate all possible combinations of the sample data and determine the proportion of these combinations that meets the chance constraints. This proportion serves as an estimate of the real confidence level of $S$. However, it is impractical due to its computational complexity of $L^{m}$, especially for large-scale problems. Therefore, we propose an exact evaluation for small-scale problems that is much faster than brute-force enumeration. For large-scale problems, we propose an accelerated Monte Carlo evaluation.